OpenStudy (anonymous):

What is the solution to the rational equation 3/x-3=2/x-2

OpenStudy (anonymous):

\[\frac{ 3 }{x-3} = \frac{ 2 }{ x-2}\]

OpenStudy (dls):

cross multiply

OpenStudy (primeralph):

or just flip both.

OpenStudy (primeralph):